OpenStudy (anonymous):

-3x-3y=3 y=-5x-17.....need to make sure im doing the right steps

OpenStudy (anonymous):

What's this? U want to solve both equations?

OpenStudy (anonymous):

you can re-write it as: -y-x=1 (divide first equation by 3) y+5x=-17 add both equation... what would you get?

OpenStudy (anonymous):

4x=-16 x=-4 put the value of x in any given equation to find y:

OpenStudy (anonymous):

I'll use: -y-x=1 then: y=-x-1=-(-4)-1=4-1=3 y=3